OpenStudy (anonymous):

I need help taking the derivative of the problem listed below, and I will show what i did. Can you please tell what I did wrong, and how I can avoid doing it again, thanks! :-) I need help taking the derivative of the problem listed below, and I will show what i did. Can you please tell what I did wrong, and how I can avoid doing it again, thanks! :-) @Mathematics

OpenStudy (anonymous):

\[\prime(x)(7x^2+8x+6)/(\sqrt{x} )\]

OpenStudy (anonymous):

show what you did :D

OpenStudy (anonymous):

In the process hahah give me 3 min

OpenStudy (anonymous):

\[F(x): x^(-.5)(7x^2+8x+6)-(x^.5) ( 14x+8)/(\sqrt{x}^2)\]

OpenStudy (anonymous):

\[F(x): (7x^\left\{ 1.5 \right\}+8x^\left\{ -.5 \right\}+6x^\left\{ -.5 \right\})-(14x^.5+8x^\left\{ -.5 \right\})/(\sqrt{x}^2)\]

OpenStudy (anonymous):

\[\prime(x):(7x^\left\{ 1.5 \right\}+6x^\left\{ -.5 \right\}-14x^.5)/(\sqrt{x}^2)\]

OpenStudy (anonymous):

I'm not sure if i can do this but I will I'm going to factor out .5 on top and in the bottom by splitting radical X squared into -> \[x^.5 \times x^.5\]

OpenStudy (anonymous):

anyone want to throw me a bone, and help me through this? Do i distribute out a X^(.5) or just the ^.5?

OpenStudy (stormfire1):

I'm unclear on the first symbol in the problem...what is that? A "1"?

OpenStudy (stormfire1):

nm...I see the second posting :) duh

OpenStudy (anonymous):

At the very beginning? Its the function symbol..shrunk. And I think, the web work will only accept the answer with the denominator intact. Not simplified

OpenStudy (stormfire1):

To be clear: the original problem is: \[x*\frac{7x^2+8x+6}{\sqrt(x)}\] ?

OpenStudy (anonymous):

instead of X * {...} its {....} / (radicalx)

OpenStudy (stormfire1):

ok, so it's just \[\frac{(7x^2+8x+6)}{\sqrt(x)}\]

OpenStudy (stormfire1):

?

OpenStudy (anonymous):

correct mate

OpenStudy (stormfire1):

ok...give me a min and let me give it a shot :)

OpenStudy (anonymous):

Thanks, i appreciate it. You will have to explain it as well, because I keep getting it wrong on the webwork..... hella frustrating haha.

OpenStudy (stormfire1):

Here's what I got for the first step:\[\frac{x^\frac{1}{2}(14x+8)-(7x^2+8x+6)(\frac{1}{2\sqrt(x)})}{x}\]

OpenStudy (anonymous):

That isn't the quotient rule Stormy :-p hahah, Its the derivative of F x G - the Derivative of G x F over G (x)^2 :-p .... X squared can be re written as X^(1/2)

OpenStudy (stormfire1):

Then I guess I won't be much help :) You said it's not the quotient rule but then you followed that statement with the definition of the quotient rule...

OpenStudy (anonymous):

u can simply simplify the term then u can easily differentiate it !!

OpenStudy (anonymous):

haahah I know, but I want to be able to replicate with confidence!

OpenStudy (anonymous):

I understood your steps, but Marvils derivative I didn't know how she got... or I couldn't follow :-X

OpenStudy (stormfire1):

I'm still trying to figure out how it's not the quotient rule.. :)

OpenStudy (stormfire1):

(loDhi-hiDlo)/lo^2 :)

OpenStudy (stormfire1):

You can also use u-substition where u = 1/sqrt(x) and v = (7x^2+8x+6)

OpenStudy (anonymous):

YOu would have to explain hahah because I'm use to quotient rule,

OpenStudy (stormfire1):

The quotient rule you posted is the same one I followed...so your comment about it totally confused me. I don't see a mistake in what I did...but if you do, please say so :) If you haven't gotten up to u-substition in your class yet, you will. It makes things a lot easier for problems like these. You can see the full steps for doing that on Wolfram's site

OpenStudy (anonymous):

Sorry, I didn't follow the rule set you were using. Now that you explained .... there is nothing wrong with. Just unusual that your dealing with a fraction with a radical in the denominator on the top... makes it seem difficult haha

OpenStudy (stormfire1):

I think if I worked it all the way through my next step would have been to convert it to a exponent. Anyway, u-substition is the way to go on this one...once you've learned it :)